Weather was calm, warm, sunny, severe clear with nary a cloud in the sky.
S7 (Reagan ranch):
S7  was still showing no data, so checked in with both Micah and Brent, and received permission to enter the site. Micah was able to open gate once I contacted him (24) from the gate.
Tried to connect to the R Pi with console cable, and with wifi; neither worked. Both Red and green LED on the Pi were solid on (Green is normally blinking).  Data had stopped recording to the raw data file at 6:27 on April 22, and we had missed ~36 hours of data.  Rebooted several times, and verified data was being recorded, also notified Gary about S7 Pi lockup.
- Reboot Pi Several time
 - Guy wires all felt tight
 - Cleaned NR01
 - NR01 ratchet strap tail was loose and blowing into sample volume, cleaned this up retied.
 
S4:
S4 was also showing all NAs. Went to site. I was able to log in with the console, and files were still being saved to the USB stick. Rebooted to get the modem back on line.
- Rebooted Pi
 - removed unterminated ribbon cable from DSM
 - Added black tape over DSM power button to reduce visibility (in highly visible location)
 - Guy wires all felt tight
 - Cleaned NR01
 - NR01 strap itself was loose, ratcheted strap about 4 clicks to get more tension.
 
Went back to base and completed weighing of 3 soil samples